OpenStudy (anonymous):

A doghouse is to be built in the shape of a right trapezoid, as shown below. What is the area of the doghouse? A right trapezoid with top side 7 feet and height 7 feet. The bottom side is extended to the right with 3 more feet. A.24.5 square feet B.59.5 square feet C.70 square feet D.147 square feet

OpenStudy (anonymous):

@aaronq @RosieF @Teddyiswatshecallsme @MMMercado @peachpi

OpenStudy (anonymous):

Will fan and medal!!

OpenStudy (anonymous):

the doghouse has the form of a right trapezoid 

OpenStudy (anonymous):

you could calculate the area separating in two areas, Area of the square + area of the triangle

OpenStudy (anonymous):

So what would that be?

OpenStudy (anonymous):

Square: \[7^{2}=49\] Triangle: \[(7*3)/2\] Square + Triangle = (21+98)/2 =59.5

OpenStudy (anonymous):

Thanks :D is it ok if I tag you on a few other questions I have on the same subject @baad1994

OpenStudy (anonymous):

it's ok man :)

OpenStudy (anonymous):

it's been 84 years.....

OpenStudy (anonymous):

im surprised ppl still mention me